OpenStudy (anonymous):

Assume f(x) = 2x + 8 and g(x) = 5. What is the value of (g o f)(3)?

OpenStudy (anonymous):

(f - g)(x) = f(x) - g(x) = -2x + 8 - 5x = 8 - 7x (f - g)(2) = 8 - 7*2 = 8 - 14 = -6

OpenStudy (oaktree):

So f(2) = -2(2)+8 = -4+8 = 4, and g(2) = 5(2) = 10. Thus, (f-g)(2) = f(2)-g(2) = 4-10 = -6.